OpenStudy (anonymous):

(a + 2b)^3 = a^3 + 6a^2b +... Use the Binomial Theorem to complete the binomial expansion.; enter only the next two terms

OpenStudy (anonymous):

\[(a+2b)^3=a^3+\left(\begin{matrix}3 \\ 1\end{matrix}\right)a^2\left( 2b \right)^1+\left(\begin{matrix}3 \\ 2\end{matrix}\right)a^1(2b)^2+(2b)^3\]\[(a+2b)^3=a^3+6a^2b+12ab^2+8b^3\]